OpenStudy (anonymous):

Find F^-1(y), the inverse of F(x) F(x) = x-10 Can somebody show me step by step how to do it

OpenStudy (harsimran_hs4):

so what F(x) does is that it gives a unique value for different values of x so let me say that for a x F(x) = x -10 = y now the inverse function gives us the values of x for which the particular y was obtained so x = 10 - y 10 -y is the required function

OpenStudy (anonymous):

so that would be the answer

OpenStudy (harsimran_hs4):

sorry a little mistake x - 10 = y x = y + 10 y+10 is the answer

OpenStudy (anonymous):

thanks

OpenStudy (harsimran_hs4):

welcome!!